Write the slope-intercept equation for the line that passes through (-3,-15) and is perpendicular to -6x+8y=3
Answer by mananth(16949)

-6 x + 8 y = 3
Find the slope of this line

8 y = 6 x + 3
Divide by 8
y = 3/4 x + 0.38
Compare this equation with y=mx+b
slope m = 0.75

The slope of a line perpendicular to the above line will be the negative reciprocal= -4/3
m1*m2=-1
m= -4/3 ,point ( -3 , -15 )
Find b by plugging the values of m & the point in
y=mx+b
-15 = 4.00 + b
b= -19
m= -4/3
The required equation is y=(-4/3)x-19
m.ananth@hotmail.ca